OpenStudy (anonymous):

C(10,3)=n! What is the value of N

OpenStudy (anonymous):

\[C(10,3)=120\]

OpenStudy (anonymous):

Howd you get it tho??

OpenStudy (anonymous):

\[\frac{10\times 9\times 8}{3\times 2}\]

OpenStudy (anonymous):

cancel first, multiply last

OpenStudy (anonymous):

@satellite73 the answer in the back of book says "5" though lol

OpenStudy (anonymous):

because it is asking for \(n\) if \(n!=120\)

OpenStudy (anonymous):

Ohhh

OpenStudy (anonymous):

\[5\times 4\times 3\times 2=120\] so \(n=5\)

OpenStudy (anonymous):

How did you get the 10X9X8 over 3X2 for your beginning equation?

OpenStudy (anonymous):

This is my first problem on the Comp Functions so I have no clue how to do them

OpenStudy (anonymous):

\[\frac{\overbrace{10\times 9\times 8}^{\text{3 terms}}}{3!}\]